Answer following questions, based on the following information on national income accounts of hypothetical economy: GDP $12,000 Gross investment 1,000 Net investment 600 Net Export 3,000 Government expenditure 1,000 Government budget balance 2,000 Net factor payment from abroad -2,000 Indirect Business Tax 2,000 A) Calculate the gross private consumption expenditure of the economy B) Calculate the Gross National Product (GNP).


1
Expert's answer
2021-09-17T09:20:13-0400

A) The gross private consumption expenditure of the economy is:

C = Y - G - I - NX = 12,000 - 600 - 3,000 - 1,000 = 7,400.

B) The Gross National Product is:

GNP = 12,000 + 2,000 = 14,000.